OpenStudy (anonymous):

8x+2y=10 Please help me solve this step by step WILL MEDAL AND FAN

OpenStudy (anonymous):

try changing a variable with 0

OpenStudy (fibonaccichick666):

So unless you are solving for a viable, you can not solve this. Are you perchance asked to solve for y?

OpenStudy (anonymous):

yes, solve the equation for Y

OpenStudy (fibonaccichick666):

ok, so can you tell me the order of operations?

OpenStudy (anonymous):

pemdas

OpenStudy (fibonaccichick666):

ok so when we are solving for a variable, we are essentially reversing the order of operations. So we do SadMep So first, is there anything being added or subtracted from our y term?

OpenStudy (anonymous):

8x

OpenStudy (fibonaccichick666):

right, and are we adding or subtracting?

OpenStudy (anonymous):

adding

OpenStudy (fibonaccichick666):

right, So how do we "undo" addition?

OpenStudy (anonymous):

by subtracting

OpenStudy (fibonaccichick666):

good, now, what will we need to subtract FROM BOTH SIDES in order to make no addition on our y term?

OpenStudy (fibonaccichick666):

then, please do that subtraction

OpenStudy (anonymous):

8x?

OpenStudy (anonymous):

8x+2y=10 -8x 2y=10-8x divide by 2 y=5-4x rewrite y=-4x+5

OpenStudy (fibonaccichick666):

yea, so make sure you subtract that from both sides

OpenStudy (fibonaccichick666):

@susie.gg98 please do not just give answers. We are here to help

OpenStudy (fibonaccichick666):

@Love_k what did you get from your subtraction?

OpenStudy (anonymous):

@FibonacciChick666 She asked for step-by-step help and thats what I did, I did not just simply "give" an answer, I showed how to get to the answer, therefore I am helping. Please do not say things that don't make sense

OpenStudy (fibonaccichick666):

What you did provides no assistance in attaining an understanding Any student could simply copy and claim it as their work, hence, just an answer.

OpenStudy (anonymous):

keep calm guys

OpenStudy (fibonaccichick666):

You also did not make a point that you were doing these things to both sides which can cause a fundamental misunderstanding. Please if you are just going to give a full answer, at least explain the concept in full instead of taking shortcuts in notation.

OpenStudy (fibonaccichick666):

My apologies @Love_k , did you still need to continue the process?

OpenStudy (anonymous):

It provides assistance by showing a student exactly how its done. I don't know about you but for me, when learning a math problem , the way you explain it does not help me whatsoever. I have to visually see the steps to get help and many others are like that. A student can copy what you are saying too making your help just an answer as well. Teachers at school give examples just like what I provided, and that doesn't mean they just simply "gave me an answer." Also I did show that I did it to both sides when I provided the statement after. Please don't be so irrelevant

OpenStudy (anonymous):

I think i know how to do it i just forgot the process of it and needed an example thank you guys!

OpenStudy (anonymous):

No problem :) I knew an example would help you

OpenStudy (fibonaccichick666):

If that were the case you would have shown a different problem, for example*. Please stop trying to pass just doing the problem and giving an answer as an example. Example * \[3x+6y=12\]First thing we need to do is get the y-term by itself. So we will subtract 3x from both sides. \[3x-3x+6y=12-3x\] Then we simplify to obtain \[6y=12-3x\] Now we need to remove the multiplication of 6 on our y-term. To do this we multiply both sides by \(\frac{1}{6}\). We then have: \[(\frac{1}{6})6y=\frac{1}{6}(12-3x)\] Using the distributive property and multiplying the left side out, we obtain: \[y=2-0.5x\] ^Above is an example problem with explanation. In future please remember to explain your work and potentially create a new question if it is just to be an "example"

OpenStudy (fibonaccichick666):

Hope the method makes sense, always try to think it out and never feel bad about writing sentences explaining what you do. As a teacher, I find it makes assisting students much easier as I see their misconceptions. Best of luck.

OpenStudy (anonymous):

LOL OMG YOU ARE FUNNY HAHAHA. Back to the original claim you made. "We are trying to help" I think I helped so..... BYE :)